OpenStudy (anonymous):

Use a decimal to find 74% of 22.

OpenStudy (anonymous):

22 x 0.74 = 16.28

OpenStudy (anonymous):

\[0.74=(0.10*7)+(0.01*4)\]\[0.10*22=2.2\]\[0.10*7=2.2*7=15.4\]\[0.01*22=0.22\]\[0.01*4=0.22*4=0.88\] \[0.74*22=15.4+0.88=16.28\]

OpenStudy (rina.r):

16.28